OpenStudy (anonymous):

Simplify the radical: 4 square root 20

jhonyy9 (jhonyy9):

any idea ?

jhonyy9 (jhonyy9):

so you know that 20=4*5

jhonyy9 (jhonyy9):

4=2^2

jhonyy9 (jhonyy9):

sqrt(5*2^2) =2sqrt5

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so 4sqrt20 = 4*2sqrt5 = 8sqrt5 hope you understand it sure

jhonyy9 (jhonyy9):

thank you for your contribution,cooperation

OpenStudy (anonymous):

sorry, i was away :/

OpenStudy (anonymous):

I dont think thats the answer though O_o